OpenStudy (anonymous):

1Which best describes the relationship between the lines? 2x – y = 5 3x – y = 5 parallel perpendicular neither same line

OpenStudy (mathmale):

Do the lines have the same, or different, slopes? If the lines have the same slope, what does that tell you about the geometric relationship between them? If two lines are perpendicular, what relationship must their slopes have?

OpenStudy (jack1):

i like the dragon yin-yang @kayeckelberg

OpenStudy (anonymous):

thank you

OpenStudy (mathmale):

@Jack1: Please let's focus on the math, and not on the dragon, OK? Also, please give @kayeckelberg a chance to answer at least some of my questions before jumping in. Thanks.

OpenStudy (anonymous):

i guess its same lines

OpenStudy (jack1):

\[y = 2x – 5\] \[y = 3x – 5\] so if \[y = mx + c\] and m is gradient and c is y intercept... what do they have in common dude? ie parallel lines have same gradient (m value) perpendicular lines have reciprocal gradient (m values) so if u had one with gradient of 1/3 and another with gradient of 3 then they'd def be perpendicular... n e ways, just hints dude, gud luck

OpenStudy (anonymous):

@jack1 what did u do?

OpenStudy (jack1):

rearranged to slope-intercept form... then compared...?

OpenStudy (anonymous):

oh yah how can i forget my bad

OpenStudy (mathmale):

@jack1: I asked you to give @kayeckelberg a chance to answer my questions. Only after he/she has responded should you be handing out the answers.

OpenStudy (jack1):

dude he or she bailed hey, look up the top: OP's gone man

OpenStudy (jack1):

besides, i didn't solve for her... only hints

OpenStudy (jack1):

aw, shes back, hiya

OpenStudy (anonymous):

I am a her! okay!

OpenStudy (anonymous):

I am very chill!

OpenStudy (jack1):

you're an abusive d1ck @mathmale , u know that? no one like's ur help, because it takes fuken forever, because ur an unhelpful pratt who couldn't teach ppl how to count, let alone complex numbers. u ask questions rather than provide similar examples, so if people don't yet have the required knowledge, ur teaching or "help" is a complete waste of time. so a great big fuk u! comin ur way buddy, how bout u try givin us another warning or ban just for helpin someone who was further confused by ur methods?

OpenStudy (anonymous):

@Jack1 its offensive language do not use it

OpenStudy (texaschic101):

back off....mathmale is a good help.....he tries to teach people how to do the work for future problems. He is good. So BACK OFF

OpenStudy (mathmale):

Thank you for your support, ChiefArnav and texaschick101. Please, could we get back to helping @kayeckelberg with her question?

OpenStudy (mathmale):

@kayeckelberg : Let's take a different approach. Please find the slope of each of the two given lines. Then please compare those slopes.

OpenStudy (anonymous):

thank you but I got on.

OpenStudy (mathmale):

You what? "got on" = ??

OpenStudy (jack1):

ya dude, looks like she solved... maybe it was the "hints" ;P